A 22.11 g sample of liquid nitrogen (N2) starts at a temperature of -288 oC. How much energy is required to heat it to its boiling point of -196 oC and boil it completely?

s = 1.04 J/goC delta Hv = 2.79 kJ/mol



1
Expert's answer
2021-04-09T01:45:56-0400

Solution:

The molar mass of N2 is 28.013 g/mol.

Hence,

Moles of N2 = (22.11 g N2) × (1 mol N2 / 28.013 g N2) = 0.789 mol N2

n(N2) = 0.789 mol


This problem can be summarized thusly:

1) Sample of liquid nitrogen (N2) is heated from -288°C to -196°C.

The heat is calculated by using the specific heat of liquid nitrogen (s = 1.04 J/g°C) and the equation:

q = m × S × ΔT

q1 = 22.11 g × 1.04 J/g°C × (-196 - (-288))°C = 2115.5 J = 2.116 kJ

q1 = 2.116 kJ


2) Liquid nitrogen is vaporized at -196°C.

The heat is calculated by multiplying the moles of liquid nitrogen by the heat of vaporization (ΔHv).

q2 = n(N2) × ΔHv = 0.789 mol × 2.79 kJ/mol = 2.201 kJ

q2 = 2.201 kJ


q = q1 + q2 = 2.116 kJ + 2.201 kJ = 4.317 kJ = 4.32 kJ

q = 4.32 kJ


Answer: 4.32 kJ of energy is required.